Find the slope and intercept of the line (x/a)+(y/b)=1.
Answer by stanbon(75887)

Find the slope and intercept of the line
(x/a)+(y/b)=1.
-----
y/b = (-x/a) + 1
-----
y = (-b/a)x + b
------
Slope = -b/a
intercept = b
------------------
Cheers,
Stan H.
============